imagine that you are planning for retirement. you expect to retire in. 35 years and you think that you can afford to save 500 dolar per month. futuremore, you believe that you can reasonably expect to earn about 8% per year without taking too much risk. how much will you have accumulated at the time that you are retire
1
Expert's answer
2020-11-09T20:38:38-0500


The term is 35 years, in months it will be:

"35\times12=420"

n=420

we will find a monthly rate

"\frac{8}{100}=0.08"

"\frac{0.08}{12}=0.00667"

r=0.00667

You can calculate the future value of a prenumerando annuity using the following formula:

"FVA=A(1+r)^n+\frac{A((1+r)^n-1)}{r}(1+r)"

"FVA=500(1+0.00667)^{420}+\frac{500((1+0.00667)^{419}-1)}{r}(1+0.00667)=1 155 729.19"